# Automatic Fused Multimodal Deep Learning for Plant Identification

### v2

@@ Abstract
Plant classification is vital for ecological conservation and agricultural productivity, enhancing our understanding of plant growth dynamics and aiding species preservation. The advent of deep learning (DL) techniques has revolutionized this field by enabling autonomous feature extraction, significantly reducing the dependence on manual expertise. However, conventional DL models often rely solely on single data sources, failing to capture the full biological diversity of plant species comprehensively. Recent research has turned to multimodal learning to overcome this limitation by integrating multiple data types, which enriches the representation of plant characteristics. This shift introduces the challenge of determining the optimal point for modality fusion. In this paper, we introduce a pioneering multimodal DL-based approach for plant classification with automatic modality fusion. Utilizing the multimodal fusion architecture search, our method integrates images from multiple plant organs—flowers, leaves, fruits, and stems—into a cohesive model. Our method achieves 82.61% accuracy on 979 classes of the PlantCLEF2015 dataset, surpassing state-of-the-art methods and outperforming late fusion by 10.33%. Through the incorporation of multimodal dropout, our approach demonstrates strong robustness to missing modalities. We validate our model against established benchmarks using standard performance metrics and McNemar’s test, further underscoring its superiority.
@@
Plant Identification , Multimodal Learning , Fusion Automation , Multimodal Fusion Architecture Search , Neural Architecture Search , Deep Learning
@@ 1 Introduction
However, typically, both traditional ML and DL models developed for plant classification tasks are constrained to a single data source, often leaf or whole plant images. From a biological standpoint, a single organ is insufficient for classification (Nhan et al., 2020), as variations in appearance can occur within the same species due to various factors, while different species may exhibit similar features. Moreover, using a whole plant image is insufficient, as different organs vary in scale, and capturing all their details in a single image is impractical (Wäldchen et al., 2018). In response to this limitation, very recent studies have delved into the application of multimodal learning techniques (de Lutio et al., 2021; Liu et al., 2016; Nhan et al., 2020; Salve et al., 2018; Hoang Trong et al., 2020; Wang et al., 2022; Zhou et al., 2021), which integrate diverse data sources to provide a comprehensive representation of phenomena. Particularly, Nhan et al. (2020) illustrate that leveraging images from multiple plant organs outperforms reliance on a single organ, in line with botanical insights. Wäldchen et al. (2018) underscore the emerging trend of multi-organ-based plant identification, indicating promising accuracy improvements due to the diverse plant viewpoints.
@@
Among these fusion strategies, late fusion emerges as the most prevalent in the observed plant classification literature, presumably due to its simplicity and adaptability (Baltrusaitis et al., 2019). However, notable work in plant classification by Nhan et al. (2020) demonstrates a compelling alternative through a hybrid approach, showcasing remarkable accuracy in a large-scale dataset. The choice of a specific strategy relies on the discretion of the model developer (Xu et al., 2021), which can introduce bias. This raises a fundamental question: what is the optimal point for modality fusion?
Contribution: In this paper, we propose a novel automatic fused multimodal DL approach for the first time in the context of plant classification. In this regard, the study first identifies the inadequacy of unimodal DL for plant classification, prompting an exploration of multimodal learning, particularly through multimodal fusion. Then, considering images of four visible plant organs—flowers, leaves, fruits, and stems—as distinct modalities, we apply the multimodal fusion architecture search algorithm (MFAS) (Perez-Rua et al., 2019) to automatically fuse these modalities. To do so, a unimodal model for each modality is first trained using the MobileNetV3Small pre-trained model. These models are then fused using the MFAS algorithm, resulting in a unified model architecture. Moreover, we evaluate the proposed model against an established baseline in the form of late fusion using an averaging strategy (Baltrusaitis et al., 2019), utilizing standard performance metrics and McNemar’s statistical test (Dietterich, 1998). The results demonstrate that our automated fusion approach enables the construction of a more effective multimodal DL model for plant identification, outperforming the baseline. Additionally, we assess the proposed model on all subsets of plant organs, revealing its robustness to missing modalities when trained with a multimodal dropout technique (Cheerla and Gevaert, 2019). The source code of our work is available online at the GitHub repository [^1] GitHub repository: https://github.com/AlfredsLapkovskis/MultimodalPlantClassifier.
@@ 2.3.1 Data Preprocessing
The original PlantCLEF2015 dataset, in its unprocessed state, lacks suitability for the intended development of our multimodal proposed model. Notably, the dataset contains images stored in varying resolutions and color models without appropriate grouping according to modalities. To rectify these issues and prepare the data and metadata for subsequent model development, essential preprocessing steps have been executed. Further details on these preprocessing steps are provided below.
@@ 4 Discussion
Our proposed multimodal DL model, which utilizes plant organ images fused through the MFAS algorithm, demonstrates high effectiveness in automating plant classification, achieving an accuracy of 82.61% and significantly surpassing the defined baseline. While a strict comparison with previous studies is not feasible due to unique aspects of our data preprocessing and methodology, Table 10 indicates that our model achieves the highest accuracy among similar research. To be more specific, for example, Ghazi et al. (2017) employed all organs from the PlantCLEF2015 dataset. Despite utilizing larger and more sophisticated classifiers, along with advanced scoring methods, they achieved an accuracy of 80.18%. Despite our architecture being relatively small (10 million parameters), the incorporation of multimodal fusion allowed us to outperform their model, highlighting the strength of multimodality and the advantages of an optimized fusion strategy of our proposed method. Similarly, de Lutio et al. (2021) sampled a large dataset comprising 56608 images of 977 species from the iNaturalist database. Utilizing images and spatio-temporal context, their multimodal model achieved accuracies of 79.12% without satellite imagery and 79.73% with it. While a direct comparison is challenging due to the use of different datasets, our approach, incorporating various organ modalities, yielded higher results. Furthermore, Lee et al. (2018), also using the PlantCLEF2015 dataset, achieved an accuracy of 68.5%. Ge et al. (2016) and Nguyen et al. (2016) extracted flower species from the PlantCLEF2015 dataset and achieved an accuracy of 52.1% and 67.45% respectively. All these results are significantly lower than ours, underscoring the effectiveness of our approach.
@@
Proposed method PlantCLEF2015 979
@@
Nguyen et al. (2016) PlantCLEF2015 967 Flower GoogleNet 67.45%
Ge et al. (2016) PlantCLEF2015 967 Flower MixDCNN 52.1%
@@
Extending the related work, we might observe some other multimodal studies achieving better results; however, this is often due to less complex experimental setups. For instance, Zhang et al. (2012) achieved an accuracy of 93.23%, but on a dataset comprising very small images. Similarly, Liu et al. (2016) achieved top-1, top-5 and top-10 accuracies of 71.8%, 91.2% and 96.4% without geographical data, and 50%, 100% and 100% with it, respecively. However, their dataset included only 50 species, each with 10 leaf and 10 flower images, collected in controlled conditions. Likewise, Salve et al. (2018) demonstrated a high GAR score, but their dataset consisted of only 60 species with 10 images each, all collected in laboratory conditions. Furthermore, Nhan et al. (2020) achieved a maximum 98.8% accuracy by utilizing all PlantCLEF2015 organs and 91.4–98.0% depending on model configurations, on only flowers, leaves, fruits, and stems, but they sampled only 50 species from the dataset, each containing all organs, thus not experiencing missing modalities as we did.
@@ 5 Conclusions
This study has addressed a critical task in agriculture and ecology: plant classification. By proposing a novel approach in this domain—a multimodal DL model utilizing four plant organs automatically fused via the MFAS algorithm, the study has demonstrated high performance, outperforming other state-of-the-art models despite the smaller size of the model. This underscores the effectiveness of multimodality and an optimal fusion strategy. Moreover, with the MD technique, the model exhibits robustness to missing modalities, even when only a single modality is available. We believe that this proposed approach opens up a promising direction in plant classification research. This highlights the need for further exploration through the development of more sophisticated algorithms capable of handling larger numbers of species, thereby unlocking the full potential of this approach. A future research could also consider incorporating a multimodal fusion of vision transformers instead of CNNs for plant classification.

### v3

@@ Abstract
Plant classification is vital for ecological conservation and agricultural productivity, enhancing our understanding of plant growth dynamics and aiding species preservation. The advent of deep learning (DL) techniques has revolutionized this field by enabling autonomous feature extraction, significantly reducing the dependence on manual expertise. However, conventional DL models often rely solely on single data sources, failing to capture the full biological diversity of plant species comprehensively. Recent research has turned to multimodal learning to overcome this limitation by integrating multiple data types, which enriches the representation of plant characteristics. This shift introduces the challenge of determining the optimal point for modality fusion. In this paper, we introduce a pioneering multimodal DL-based approach for plant classification with automatic modality fusion. Utilizing the multimodal fusion architecture search, our method integrates images from multiple plant organs—flowers, leaves, fruits, and stems—into a cohesive model. To address the lack of multimodal datasets, we contributed Multimodal-PlantCLEF, a restructured version of the PlantCLEF2015 dataset tailored for multimodal tasks. Our method achieves 82.61% accuracy on 979 classes of Multimodal-PlantCLEF, surpassing state-of-the-art methods and outperforming late fusion by 10.33%. Through the incorporation of multimodal dropout, our approach demonstrates strong robustness to missing modalities. We validate our model against established benchmarks using standard performance metrics and McNemar’s test, further underscoring its superiority.
@@
Plant Identification , Multimodal Learning , Fusion Automation , Multimodal Fusion Architecture Search , Neural Architecture Search , Multimodal Dataset
@@ 1 Introduction
However, typically, both traditional ML and DL models developed for plant classification tasks are constrained to a single data source, often leaf or whole plant images. From a biological standpoint, a single organ is insufficient for classification (Nhan et al., 2020), as variations in appearance can occur within the same species due to various factors, while different species may exhibit similar features. Moreover, using a whole plant image is insufficient, as different organs vary in scale, and capturing all their details in a single image is impractical (Wäldchen et al., 2018). In response to this limitation, very recent studies have delved into the application of multimodal learning techniques (de Lutio et al., 2021; Liu et al., 2016; Nhan et al., 2020; Salve et al., 2018; Hoang Trong et al., 2020; Wang et al., 2022; Zhou et al., 2021), which integrate diverse data sources to provide a comprehensive representation of phenomena. Particularly, Nhan et al. (2020) illustrate that leveraging images from multiple plant organs outperforms reliance on a single organ, in line with botanical insights (Goëau et al., 2015). Wäldchen et al. (2018) underscore the emerging trend of multi-organ-based plant identification, indicating promising accuracy improvements due to the diverse plant viewpoints.
@@
Among these fusion strategies, late fusion emerges as the most prevalent in the observed plant classification literature, presumably due to its simplicity and adaptability (Baltrusaitis et al., 2019). However, the choice of a specific strategy relies on the discretion of the model developer (Xu et al., 2021), which can introduce bias and lead to a suboptimal architecture.
To address these issues, we propose an automatic multimodal fusion approach utilizing images from four distinct plant organs—flowers, leaves, fruits, and stems. Following Nhan et al. (2020), we refer to these organs as modalities. While all organs are represented as RGB images, each encapsulates a unique set of biological features, reflecting the fundamental property of multimodality— complementarity (Lahat et al., 2015). Furthermore, unlike multi-view methods (Seeland and Mäder, 2021), our model explicitly requires distinct plant organs as inputs. Additionally, unlike certain multi-organ methods (Lee et al., 2018), our model has a fixed set of inputs, with each input corresponding exclusively to a specific organ. Thus, we suggest the term multimodality aligns more accurately with our approach.
That said, existing plant classification datasets are predominantly designed for unimodal tasks, which poses a significant challenge for developing and evaluating multimodal approaches. To address this limitation, we introduce a data preprocessing pipeline that transforms an existing unimodal dataset, namely PlantCLEF2015 (Joly et al., 2015), into a multimodal dataset comprising combinations of plant organ images. This dataset, which we call Multimodal-PlantCLEF, supports the development of models with a fixed number of inputs, each corresponding to a specific plant organ.
Our contribution is fourfold:
  1. 1.
    We propose a novel data preprocessing approach to convert a unimodal plant classification dataset into a multimodal one and apply it to transform PlantCLEF2015 into Multimodal-PlantCLEF for use in this study.
  2. 2.
    We propose a novel automatic fused multimodal DL approach for the first time in the context of plant classification. To do so, we first train a unimodal model for each modality using the MobileNetV3Small pre-trained model. Then, we apply a modified multimodal fusion architecture search algorithm (MFAS) (Perez-Rua et al., 2019) to automatically fuse these unimodal models.
  3. 3.
    We evaluate the proposed model against an established baseline in the form of late fusion using an averaging strategy (Baltrusaitis et al., 2019), utilizing standard performance metrics and McNemar’s statistical test (Dietterich, 1998). The results demonstrate that our automated fusion approach enables the construction of a more effective multimodal DL model for plant identification, outperforming the baseline.
  4. 4.
    Additionally, we assess the proposed model on all subsets of plant organs, revealing its robustness to missing modalities when trained with a multimodal dropout technique (Cheerla and Gevaert, 2019).
The source code of our work is available online at the GitHub repository [^1] GitHub repository: https://github.com/AlfredsLapkovskis/MultimodalPlantClassifier.
@@ 2.3.1 Data Preprocessing
The original PlantCLEF2015 dataset and its associated evaluation protocols are not suitable for the development of our proposed multimodal model. These protocols are designed for models that accept either a single image or an observation comprising an arbitrary number of images from various plant organs. In contrast, our model requires a fixed input format consisting of four specific plant organ images. This fundamental mismatch necessitates a custom strategy to reorganize the original dataset into fixed combinations of plant organ images, which inherently precludes the use of standard PlantCLEF2015 evaluation protocols. Given these constraints, modifying the dataset is both a justified and essential step. It is important to note that all multimodal approaches are dealing with this issue in plant identification. For more information, see Section 4.4. In the following, we detail our preprocessing pipeline for transforming PlantCLEF2015 into Multimodal-PlantCLEF.
@@ 4 Discussion
4.1 Accuracy
Our proposed multimodal DL model, which utilizes plant organ images fused through the MFAS algorithm, demonstrates high effectiveness in automating plant classification, achieving an accuracy of 82.61% and significantly surpassing the defined baseline. Table 10 indicates that our model achieves the highest accuracy among similar research.
To be more specific, for example, Ghazi et al. (2017) employed all organs from the PlantCLEF2015 dataset. Along with advanced scoring methods, they achieved an accuracy of 80.18%. Similarly, de Lutio et al. (2021) sampled a large dataset comprising 56608 high-quality images of 977 species from the iNaturalist database. Utilizing images and spatio-temporal context, their multimodal model achieved accuracies of 79.12% without satellite imagery and 79.73% with it. Furthermore, Lee et al. (2018), also using the PlantCLEF2015 dataset, achieved an accuracy of 68.5%. Ge et al. (2016) and Nguyen et al. (2016) extracted flower species from the PlantCLEF2015 dataset and achieved an accuracy of 52.1% and 67.45% respectively. These results are significantly lower than ours, underscoring the effectiveness of our approach.
Extending the related work, we might observe some other multimodal studies achieving better results; however, this is often due to less complex experimental setups. For instance, Zhang et al. (2012) achieved an accuracy of 93.23%, but on a dataset comprising very small images. Similarly, Liu et al. (2016) achieved top-1, top-5 and top-10 accuracies of 71.8%, 91.2% and 96.4% without geographical data, and 50%, 100% and 100% with it, respecively. However, their dataset included only 50 species, each with 10 leaf and 10 flower images, collected in controlled conditions. Likewise, Salve et al. (2018) demonstrated a high GAR score, but their dataset consisted of only 60 species with 10 images each, all collected in laboratory conditions. Furthermore, Nhan et al. (2020) achieved a maximum 98.8% accuracy by utilizing all PlantCLEF2015 organs and 91.4–98.0% depending on model configurations, on only flowers, leaves, fruits, and stems, but they sampled only 50 species from the dataset, each containing all organs, thus not experiencing missing modalities as we did. Seeland and Mäder (2021) achieved a maximum accuracy of 94.25% on the PlantCLEF2015 dataset; however, their results were based on a very small selection of images and classes.
@@ 4.1 Accuracy
4.2 Computational Cost
While our approach involves much one-time computational cost during training to identify an optimal architecture, the resulting model is highly efficient during inference. With only 10 million parameters, our model is lightweight and easily deployable to mobile devices. This parameter count could be further reduced by employing fewer fusion layers, if needed.
In comparison, de Lutio et al. (2021) utilize ResNet-50 with 25 million parameters, Lee et al. (2018) design a custom architecture exceeding 50 million parameters, and Ghazi et al. (2017) employs VGG and GoogleNet, which contain 138 million and 6.8 million parameters, respectively. Contrary, our model is significantly smaller, more efficient, and suitable for deployment on energy-constrained devices. This underscores the importance of multimodality and an optimal fusion strategy.
@@ 4.2 Computational Cost
Proposed method Multimodal-PlantCLEF 979
@@
Nguyen et al. (2016) Modified PlantCLEF2015 967 Flower GoogleNet 67.45%
Ge et al. (2016) Modified PlantCLEF2015 967 Flower MixDCNN 52.1%
@@
4.3 Search Efficiency
The MFAS algorithm incorporates several optimizations—including surrogate modeling, weight sharing, reduced model complexity, and limiting training to a small number of epochs per configuration—that enable efficient exploration of a large number of architectures. The primary bottleneck in the original procedure occurs during the first iteration, where all generated configurations must be trained to provide data for the surrogate model. To address this, we divided the initial architectures into 20 batches and trained them in parallel.
This approach allowed the algorithm to evaluate up to $2592+50\times 19=3542$ unique architectures within a time frame roughly equivalent to training a single multimodal model for $((2592/20)+50\times 19)\times 2=2159.2$ epochs. Even if the first iteration were performed sequentially, the process would involve $7084$ epochs—a feasible one-time computational cost for identifying an optimal architecture. Moreover, these requirements could be further reduced by decreasing the number of layers considered for fusion, the number of iterations performed or the number of architectures explored.
In comparison, manual search for an optimal configuration would require to investigate a search space of size $4.51\times 10^{13}$, making the MFAS algorithm not only efficient but also practical for large-scale multimodal learning tasks.
In this study, we employed MFAS as it is a well-established multimodal NAS algorithm that satisfied our efficiency requirements when compared to alternative methods (e.g., (Xu et al., 2021)). By being the first to apply such methods to plant classification, we establish a benchmark for future research in this area. Future studies in plant classification can investigate other multimodal NAS approaches to further enhance training efficiency.
4.4 Limitations
In this study, we modify the PlantCLEF2015 dataset, which is a common practice in the literature (Seeland and Mäder, 2021; Nhan et al., 2020; Nguyen et al., 2016; Ge et al., 2016). This modification was necessary due to the lack of multimodal datasets and the inherent limitations of the original PlantCLEF2015 dataset, which is unsuitable for tasks requiring fixed inputs of specific plant organs. To address this gap and promote comparability and reproducibility in future research, we contribute a preprocessing pipeline that transforms PlantCLEF2015 into Multimodal-PlantCLEF.
@@ 5 Conclusions
This study has addressed a critical task in agriculture and ecology: plant classification. By proposing a novel approach in this domain—a multimodal DL model utilizing four plant organs automatically fused via the MFAS algorithm, the study has demonstrated high performance, outperforming other state-of-the-art models despite the smaller size of the model. This underscores the effectiveness of multimodality and an optimal fusion strategy. Moreover, with the MD technique, the model exhibits robustness to missing modalities, even when only a single modality is available. Additionally, we contributed Multimodal-PlantCLEF, a restructured version of the PlantCLEF2015 dataset tailored for multimodal tasks, to support further research in this area. We believe that this proposed approach opens up a promising direction in plant classification research. This highlights the need for further exploration through the development of more sophisticated algorithms capable of handling larger numbers of species, thereby unlocking the full potential of this approach. A future research could also consider incorporating a multimodal fusion of vision transformers instead of CNNs for plant classification.